OpenStudy (anonymous):

Solve using the substitution method. 5x − 7y = −16, 2x + 8y = 26

OpenStudy (mertsj):

Divide the second equation by 2 and solve it for x You will get x=13-4y Put that into the first equation in place ofx

OpenStudy (anonymous):

i would use the 2nd equation to substitute because the numbers are more simple. get down to x =( ) so 2x +8y = 26 2x = 26 -8y x = (26 - 8y)/2 x = 13 -4y now you can replace this x into the other equation 5x - 7y = -16 replace the x 5(13 - 4y) - 7y = -16 65 - 20y - 7y = -16 solve for y -27y = -71 y = 3 replace y=3 into any equation to find x 2x + 8y = 26 2x + 8(3) = 26 2x + 24 = 26 2x = 2 x =1 y = 3

OpenStudy (anonymous):

Ok mertsj I am still confused

OpenStudy (mertsj):

Can you divide the second equation by 2?

OpenStudy (mertsj):